Sol Summary – February 29th

Crew 222 Sol Summary Report 29-02-2020
Sol: 13
Summary Title: Martian friendship
Author’s name: Benjamin Auzou
Mission Status: Crew transition with Crew 223

Sol Activity Summary: We woke up at 5:30 am and enjoyed the sunrise from
Hab Ridge. Then we headed to Loa to get the HabCar back. On our way
back, we met the crew 223 in Hanksville. That was an intensoe and joyful
moment to see our friends, seven people we prepared this mission with !
Then we spend the afternoon getting water for the tanks and explaining
each member the keys of his role. We will enjoy all a burger together
tonight, our first Earth food, and their last !

Look Ahead Plan: Tomorrow we will leave the campus very early, and we
wish a very good mission to our friends of crew 223.

Anomalies in work: none
Weather: cloudy and warm
Crew Physical Status: all members safe and sound
EVA: none
Reports to be filed: Journalist report, Operations report, GreenHab
Report, Astronomy Report, Photos Report, Photo
of the day, Mission Checklist
Support Requested: nothing to report

GreenHab Report – February 29th

  

Crew 222 GreenHab Report - 29/02/2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Both heating and ambiant, door open during
the afternoon

Average temperatures: NA

Hours of supplemental light: 19h30 PM to 00h

Daily water usage for crops: Non known

Daily water usage for research and/or other purposes: 0 Liters

Water in blue Tank: 0 gallons

Time(s) of watering crops: 2

Changes to crops: NA

Narrative:Today, Dr. Ruppert took care of the GreenHab and i don't have
all the informations

Harvest:

Support/supplies needed: NA

Journalist Report – February 29th

Crew 222 Journalist Report 29Feb2020
Author: Marie Bochard, Crew Journalist
Sol 13
Title: A taste of home

We woke up very early this morning to see the sunrise. At 6am, we were
all outside climbing on the Ridge wall in the pale light of our
flashlights. The view from above was stunning: the sun rose over the
small hills in the distance, illuminating the mesa and the snowy
mountains minutes after minutes. Finally, the sun rays shone through the
atmosphere… What a way to end our stay here.

At 1pm, we met with Crew 223 at the Rock Shop. It was so crazy seeing
new faces after two weeks in isolation! Especially since we know them
from back home, it feels like they brought a taste of home with them. We
cooked lunch and shared our experience here with the new crew. We’ve
been preparing this mission for more than a year, so being here together
is long awaited! We were really excited to tell them all about our
mission and transmit our newly acquired experience. We had so much to share.

Before the sun was setting, we all got together outside to take the crew
change photos. We are leaving early tomorrow morning, so we needed to
finish all our transmission of knowledge before we went out to eat. Once
that was all done, we headed out to have our first meal and their last
meal outside the station: the iconic “last burger” in Hanksville.

It has been a pleasure writing everyday about our life in the station.
It’s been an incredible experience for all of us, and we would love to
spend the next two weeks here again (granted we can take a shower
before…). I hope I could transmit my passion for this mission through
these reports, and I hope they made you aspire to being a marsonaut,
like it inspired me!

Operations Report – February 29th

  


 Crew 222 Operations Report 02/29/2020

SOL: 13

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Friday generator:
Off at 5:00am
On at 9:00pm SOC was 85%

Saturday generator:
Off at 5:30am
On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 82%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 43 gallons

Water Meter: 149628.91 units

Water (static tank) – 400 gallons

Static to Loft Pump used – yes

Water in GreenHab – 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no
Hours: 146.7
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: no
Hours: 132.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: no
Hours: 138.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes
Reason for use: Run during the engineer EVA for the battery.
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: Oil checked. The Honda and the ATV number 1's oil light
lit up.

HabCar used and why, where? The HabCar is at thr MDRS

CrewCar used and why, where? The CrewCar is at the MDRS

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Operations Report – Feb 28th

Crew 222 Operations Report 28FEB2020

SOL: 12

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Thursday generator:

Off at 5:30am

On at 9:00pm SOC was 83%

Friday generator:

Off at 5:00am

On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 83%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149590.24 units

Water (static tank) – 176 gallons

Static to Loft Pump used – yes

Water in GreenHab – 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: no

Hours: 146.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Spirit rover: no

Hours: 132.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Curiosity rover used: no

Hours: 138.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)

Reason for use: Run during the engineer EVA for the battery and for the

EVA this morning.

Oil Added? no

ATV Fuel Used: 20% of the tank.

Hours the ATVs were Used today: 45 minutes

Notes on ATVs: Oil checked. The Honda and the ATV number 1’s oil light lit up. At the end of the EVA, the red one couldn’t change gear. I stopped and it was stuck in 5th gear. We had to push it to change gear. When we finally put it in neutral, we started it again and it worked well until the end (the gears were working fine).

HabCar used and why, where? The HabCar is now in the shop awaiting repairs.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA. The suit number 7 was working well at the beginning but stopped working in the middle of the EVA. We fixed this in the RAM during this afternoon and now it’s working well.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Fixing of the suit number 7. It’s working well now. Finished the spacesuit experiment and all working well now.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Journalist Report – Feb 28th

Crew 222 Journalist Report 28Feb2020

Author: Marie Bochard, Crew Journalist

Sol 12

Title: First breath of fresh air

To enjoy fully our last simulation day, we started the day off with a long EVA. Nina, Eishi, Valentin and I each took an ATV and headed North to the Moons and Copernicus Highway (a dust path, as the name doesn’t suggest). The landscape was incredible. Every step we took, the ground was a different color. It really felt like we were exploring an alien planet. We found plenty of fossils of oysters, proof that there was plenty of water here a long time ago. There were glass-like shards on the floor, possibly made by lightning hitting the sand… that’s our guess. We surely walked where no one has every walked before, and it

created a very unique mix between feeling special and feeling small. After a couple hours of exploring, we headed back on the ATVs and everyone enjoyed the ride to the Hab.

The rest of the day was mostly spent cleaning. We collected our belongings and started packing.

Then it was over. We opened the Hab door at 5:30pm, finally breathing fresh air after two weeks here. How lucky are we Earthlings to be able to breathe outside and to walk freely without having to wear a pressurized suit!

This mission has gone by tremendously fast. It feels like we just got here, and yet it’s been ages since we thought of the world outside. Everyone enjoyed the stay here, and we feel closer than ever. Missions like this one can create ever-lasting bonds, sometimes more than years can achieve. It will be forged in everyone’s memory for the rest of our lives, I can guarantee that.

Operations Report – Feb 27th

Crew 222 Operations Report 27FEB2020

SOL: 11

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Wednesday generator:

Off at 6:00am

On at 9:00pm SOC was 85%

Thursday generator:

Off at 5:30am

On at 9:00pm SOC was 83%

Diesel Reading – 72%

Station Propane Reading – 84%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 200 gallons

Water Meter: 149548.32 units

Water (static tank) – 218 gallons

Static to Loft Pump used – yes

Water in GreenHab – 35.5 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes

Hours: 146.7

Beginning charge: 100%

Ending charge: 89%

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Spirit rover: no

Hours: 132.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Curiosity rover used: no

Hours: 138.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1 and 2)

Reason for use: Run during the engineer EVA for the battery and for the

EVA this morning.

Oil Added? no

ATV Fuel Used: 20% of the tank.

Hours the ATVs were Used today: 30 minutes

Notes on ATVs: Oil checked. The Honda needs an oil change. The four ATVs

were operational this afternon.

HabCar used and why, where? The habcar is now in the shop awaiting repairs.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA.

The suit number 7 was working well at the beginning but two hours after the beginning of the EVA the fan was stopping at random moments but was restarting after. The fan never stopped totally but wasn’t working optimally.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

GreenHab Report – Feb 28th

Crew 222 GreenHab Report – 28FEB2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Both heating and ambient, door open during the afternoon

Average temperatures: 21°C

Hours of supplemental light: 19h30 PM to 00h

Daily water usage for crops: 10,6 ( 6 from the static tank since the Blue tank is empty)

Daily water usage for research and/or other purposes: 0 Liters

Water in blue Tank: 0 gallons

Time(s) of watering crops: 2

Changes to crops: NA

Narrative:This morning, I went to the Greenhab to water plants, I uses water of the blue tank until it was empty so I started using water from the static tank.

Tonight i watered again plants, I also harvested some herbs for our tonight pie.

Harvest:

Oregano: 10g

Chives:8g

Sage:2g

Basil:13g

Raddishes:7g

Cilantro:5g

Support/supplies needed: There is no water left in the blue tank and I’ll have to use water from the static tank again tomorrow.